A mother is now 3 times as old as the daughter. twelve years ago she was six times as old as her daughter, how old is the daughter and mother
Answer by JulietG(1812)

M = 3D
M-12 = 6(D-12)
Substitute the value of M from the first equation into the second.
(3d-12) = 6(D-12)
Distribute
3D-12 = 6D-72
Subtract 3D from each side
-12 = 3D-72
Add 72 to each side
60 = 3D
Divide each side by 3
20 = D
.
If the daughter is 20, then the mother is 60.
Twelve years ago, the daughter would have been 8 and the mother would have been 48.